Continuing preparations for the two upcoming Russian Orlan EVAs, CDR Volkov & FE-1 Kononenko today focused on the second spacewalk, EVA-20, starting out with a three-hour review of familiarization material, watching a training video, and studying the preliminary EVA timeline (details, see below).
Later, Volkov gathered equipment and tools for EVA-20, going by an uplinked search list and supported by ground specialist tagup. (Besides two well-stocked tool kits, EVA-20 equipment includes the new VSPLESK ('Burst) science payload, a container for the BIORISK-MSN payload and a mounting platform.)
After configuring STTS communication/telemetry links for their stay in the DC1 Docking Compartment, Volkov & Kononenko then -
Readied hardware and tools for the first spacewalk, EVA-20A,
Reviewed staging procedures & photography requirements for the docking target installation, 'Yakor' foot restraint transfer, and inspection of mounting holes for the Kurs antenna,
Performed a pressure check on the supplementary BNP portable air repress bottle in the SM PkhO (Service Module Transfer Compartment),
Updated the airlock depress/repress cue cards in the DC1 'Pirs', and
Restored nominal STTS comm settings in the DC1.
FE-2 Chamitoff continued checking out JEM (Japanese Experiment Module) 'Kibo' payload systems, today working on the JCGSE (JEM Common Gas Support Equipment), with the Ar (Argon) gas supply from the Ar GBUs (Gas Bottle Units) opened to take pressure readings for leak checking. (Checkout activities included accessing and setting the JEM HRMS (High Rate Data Multiplexer & Switcher), after configuring cable connections to the HRMS patch panel. After the checkout, the Ar gas supply was stopped and the HRMS access door closed again.)
Chamitoff also conducted 'Week 11' sampling of potable water for chemical and microbial analysis from the SVO-ZV tap and two SRV-K taps, the latter after preliminary heating of the water (three heating cycles) and flushing. (Gregory collected three 225 mL samples (for inflight microbial analysis) and two 750 mL samples (for postflight chemical analysis) from each of three ports (SRV-K hot, SRV-K warm, SVO-ZV) for return on STS-126/ULF2.)
Later, the FE-2 performed the in-flight microbial analysis using the WMK (Water Microbiology Kit) and its MCDs (Microbial Capture Devices) with Coliform Detection Bags. The small amounts of flush and leftover water in the 50 mL waste water & micro sample in-flight analysis bags were reclaimed for use. (Coliform bacteria are the commonly-used bacterial indicator of sanitary quality of foods and water. They are defined as rod-shaped 'gram-negative' non-spore forming organisms that ferment lactose with the production of acid and gas when incubated at 35-37 degC.)
Gregory also conducted the weekly 10-min. CWC (Contingency Water Container) audit as part of on-going WRM (Water Recovery & Management) assessment of onboard water supplies. (Updated 'cue cards' based on the crew's water calldowns are sent up every other week. The current card (17-0002O) lists 36 CWCs (~1365.2 L total) for the four types of water identified on board: technical water (650.6 L, for Elektron, flushing, hygiene, incl. 509.4 L non-usable water because of Wautersia bacteria), potable water (706.7 L, incl. 260.6 L currently on hold), condensate water (0 L), waste/EMU dump and other (7.9 L). Wautersia bacteria are typical water-borne microorganisms that have been seen previously in ISS water sources. These isolates pose no threat to human health.)
Oleg Kononenko took the periodic readings of potentially harmful atmospheric contaminants in the SM, using the CMS (Countermeasure System), a component of the GANK-4M Real-Time Harmful Contaminant Gas Analyzer suite, which uses preprogrammed microchips to measure H2CO (Formaldehyde, methanal), CO (Carbon Monoxide) and NH3 (Ammonia), taking one measurement per microchip. Today's measurements also looked for O3 (Ozone), C6H6 (Benzene) and NO2 (Nitrogen Dioxide), using special chips. (CMS is a subsystem of the Russian SKDS Pressure Control & Atmosphere Monitoring System.)
Later tonight, Sergey Volkov is scheduled to perform monthly maintenance on the deactivated Russian IK0501 GA (Gas Analyzer) of the SOGS Pressure Control & Atmospheric Monitoring System by replacing its CO2 filter assembly (BF) with a new unit from FGB stowage (done last: 5/22), then reactivating the unit.
Kononenko has another 30 min. set aside to continue relocating and loading common trash into the ATV1 (Automated Transfer Vehicle) 'Jules Verne', based on an uplinked checklist for discarded cargo.
The FE-1 conducted today's routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables.
The CDR completed the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).
Greg rearranged samples in the MELFI (Minus Eighty Degree Laboratory Freezer for the ISS), relocating the latest blood and urine samples from their current box module in a dewar to another box module in a different dewar. (Sample vials (separated in bags) were transferred from Dewar 3 to Dewar 2 and from Dewar 1 to Dewar 2.)
The three crewmembers conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), RED resistive exercise device (CDR, FE-1, FE-2), and VELO bike with bungee cord load trainer (CDR, FE-1). Later tonight, Kononenko transfers the exercise data file to the MEC laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).
Progress M-64/29P Line Purging: Two Progress 29P prop line purges are scheduled tonight at ~6:50pm and ~8:30pm. ISS attitude control authority will be handed over from US CMG momentum management to Russian MCS (Motion Control System) thrusters at 6:35pm for the subsequent maneuvers to the line purge attitude (yaw/pitch/roll: 0.0/340.0/0.0). Return to LVLH TEA (Local Vertical/Local Horizontal Torque Equilibrium Attitude) follows at ~9:00pm and control handover back to US momentum management at ~10:40pm. For the propellant venting, the protective Lab science window shutters will be closed by Gregory at ~6:00pm and not opened until about 1:40am.
EVA-20 Timeline Preview (preliminary): The Orlan EVA-20 by Volkov (EV1) & Kononenko (EV2) on 7/15 is scheduled to begin at ~1:14pm EDT (DC1 EV hatch open), to last an estimated 5 hrs 30 min, i.e., concluding at approximately 6:44pm. Part of the EVA will be supported by the DC1-based Strela 1 crane, operated via hand crank. Main objectives of the Russian spacewalk are -
Closeout ops on Soyuz TMA-12 Plane I inspection/pyrobolt removal (if not completed during EVA-20A);
Installation of a new docking target on SM PkhO (Transfer Compartment), Plane IV, for MLM (if not completed during EVA-20A);
Post-installation photography of the new docking target;
Inspection of mounting holes for an adapter of a Kurs antenna (4AO-VKA) on PkhO-RO (SM Working Compartment, small diameter section);
Transferring one 'Yakor' foot restraint (of two) from the DC1 EVA ladder to the SM and installing it in an attachment socket at a PkhO handrail (two Yakors were installed on DC1 ladder during EVA-17A on 2/22/07);
Installing the VSPLESK ('Burst') science payload on the SM RO (large diameter section); and
Removing the BIORISK-MSN (BIO-2) experiment container 1 from the DC1 for return to the station (BIORISK-MSN #1 was installed with two other BIO-2 containers during EVA-19 on 6/6/07).
Uncrewed Station Ops Preview: One of the contingencies associated with the Orlan EVA-20A next week (and EVA-20 on 7/15) could require uncrewed station operation for some time. Most of the US Segment preparations for this eventuality will be done by Gregory Chamitoff next Wednesday (7/9), such as transferring selected hardware to the RS (Russian Segment), setting up a PCS (Portable Computer System) laptop in the FGB as backup to the PCS in the RS, powering down ham/amateur radio equipment, reconfiguring some LAN software (NetMeeting, KFX), powering down the COL PWS (Columbus Orbital Laboratory Portable Workstation) laptop, and closing selected hatches. On 7/10, the IATCS (Internal Thermal Control System) will be configured as usual for uncrewed ops, some racks will be jumpered to the LTL (Low Temperature Loop) in case they need cooling, and some remaining hatches will be closed, before Gregory moves to the Soyuz Descent Module (SA) and closes the hatch between it and the Orbital Module (BO).
CEO (Crew Earth Observations) photo targets uplinked for today were Mt. Etna (this famous European Volcano is located in the eastern part of the island of Sicily. As ISS approached the island from the NW. Greg was to begin looking for the volcano to the E and try for detailed views of the eastern slopes of the volcano between the summit and the coast), Red River Basin, TX (the Red River is a major tributary of the Mississippi River that also forms the border between the states of Texas and Oklahoma. The basin area north of the Dallas-Fort Worth area is primarily an agricultural region that is changing under increasing urban development. As the station approached this area from the NW near midday, Greg was to look for the W-to-E meandering flow of the river and map the area west of the man-made Lake Texoma), and Santa Barbara Coast, California (the coastal area near Santa Barbara, California has been designated as a Long Term Ecological Research (LTER) site. On this pass as ISS approached the coast from the NW, looking just left of track and trying for a detailed mapping strip of the coastal range known as the Santa Inez Mountains from Point Conception to just north of the city of Santa Barbara).